Sam tosses a ball horizontally off a footbridge at 3.1 m/s. How much time passes after he releases it until its speed doubles?
1
Expert's answer
2010-12-22T03:26:42-0500
The horizontal projection is constant while the vertical is changing according to the equation
Vy=V0y+gt
(we chose the downward direction as positive).
Vx = 3.1 m/s
V0y = 0;
V0 = 3.1 m/s
V = √(Vx2 + Vy2) = √(3.12 +100t2)
V2/V02 = 4
(3.12 +100t2)/ 3.12 = 4
t =√(3.12*3 / 100) = 0.54 s